### v2

# DBCopilot: Scaling Natural Language Querying to Massive Databases

Abstract.
NL2SQL simplifies database interactions by enabling non-experts to convert natural language (NL) questions into Structured Query Language (SQL) queries. While recent advances in large language models (LLMs) have improved the zero/few-shot NL2SQL paradigm, existing methods face scalability challenges when dealing with massive databases. This paper introduces DBCopilot, a framework that addresses these challenges by employing a compact and flexible copilot model for routing over massive databases. Specifically, DBCopilot decouples schema-agnostic NL2SQL into domain-specific schema routing and generic SQL generation. This framework utilizes a lightweight differentiable search index to construct semantic mappings for massive database schemas and navigates natural language questions to their target databases and tables in a relation-aware, end-to-end manner. The routed schemas and questions are then fed into LLMs for effective SQL generation. Furthermore, DBCopilot introduces a reverse schema-to-question generation paradigm that can automatically learn and adapt the router over massive databases without manual intervention. Experimental results demonstrate that DBCopilot is a scalable and effective solution for schema-agnostic NL2SQL, providing a significant advance in handling large-scale schemas of real-world scenarios.
Text-to-SQL, LLM, Generative Retrieval, Schema Routing
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Information systems Structured Query Languageccs: Information systems Top-k retrieval in databasesccs: Information systems Question answering
Figure 1. The DBCopilot framework. DBCopilot decouples NL querying over massive databases into schema routing and SQL generation through LLM-Copilot collaboration.
1. Introduction
Natural language to SQL (NL2SQL) aims to convert natural language (NL) questions into Structured Query Language (SQL) queries, making it convenient for non-experts to query and analyze databases. This task plays a critical role in data analysis, business intelligence, and other database interaction applications, thus has attracted significant attention (Androutsopoulos et al., 1995; Affolter et al., 2019; Qin et al., 2022; Katsogiannis-Meimarakis and Koutrika, 2023; Fu et al., 2023; Gu et al., 2023a). The recent emergence of large language models (LLMs) brings a new zero/few-shot NL2SQL paradigm that directly transforms SQL queries via LLMs with schema-aware prompts, showing great promise (Rajkumar et al., 2022; Liu et al., 2023a; Pourreza and Rafiei, 2023; Gu et al., 2023b; Gao et al., 2023).
However, existing work in this direction typically focuses only on querying a single database with a handful of tables, neglecting the real-world requirement to query over massive databases and data warehouses (Pride, 2023; Kothyari et al., 2023; Sen et al., 2020). In this case, data consumers need to invest significant effort to manually identify and select the appropriate database and relevant tables. Therefore, it becomes imperative to develop a mechanism that abstracts non-expert users from the intricacies of data representation and enables them to query over massive databases with schema-agnostic NL questions, i.e., without understanding of large-scale schemas. The core of this process is to construct semantic mappings from diverse NL questions to massive database model elements, thus automatically identifying the target database and filtering for a minimal amount of relevant tables.
The above scenarios pose significant challenges to current LLM-based NL2SQL methods. Due to token limits and inference costs, it is infeasible to feed hundreds/thousands of schemas into LLMs for each question. Furthermore, current LLMs still struggle to effectively leverage information present in long contexts (Liu et al., 2023b; Xu et al., 2023). To address this scalability problem, existing solutions can be categorized into retrieval-based and fine-tuning based methods. Given an NL question, retrieval-based methods identify the corresponding table schemas using information retrieval techniques (Wikipedia contributors, 2023a; Karpukhin et al., 2020). Unfortunately, by relying on the co-occurrence of words and representations between question and schema, retrieval-based methods suffer from vocabulary mismatch and the diversity of NL expressions. In addition, these ad hoc retrieval methods typically treat documents as independent objects, neglecting the relationships between databases and tables. Fine-tuning based methods inject schema knowledge into the parameters of LLMs by fine-tuning with schema descriptions or (question, schema) pairs. Unfortunately, fine-tuning LLMs with billions of parameters to keep up with evolving database schemas is resource-intensive and sometimes impractical. Besides, it remains challenging to effectively inject new knowledge into LLMs through fine-tuning (Ovadia et al., 2023). As a result, how to extend the scalability of LLM-based NL2SQL methods and apply them to massive databases is a major challenge in this field.
In this paper, we present DBCopilot, a framework for effectively scaling natural language querying to massive databases. The main idea behind DBCopilot is to introduce a compact and flexible copilot model to serve as a router across massive databases, and to leverage the efficient NL2SQL capability of LLMs to generate final SQL queries. To this end, as shown in Figure 1, DBCopilot decouples schema-agnostic NL2SQL into domain-specific schema routing and generic SQL generation. Specifically, DBCopilot utilizes a lightweight differentiable search index (DSI) (Tay et al., 2022) to navigate an NL language question to its corresponding database and tables. Then the routed schemas and questions are fed into LLMs to generate SQL queries by leveraging their strong language understanding and SQL generation capabilities. In this way, DBCopilot can easily integrate with advanced LLM-based NL2SQL solutions and rapidly and cost-effectively scale to massive databases.
For an effective implementation of DBCopilot, three major challenges need to be addressed. (1) Schema routing is not a task of retrieving independent tables because complex queries require information across multiple tables, and there are inherent relationships between databases and tables. To resolve this challenge, we model schema routing as a sequence-to-sequence (Seq2Seq) DSI. By proposing schema graph construction, DFS serialization, and constrained decoding, we can jointly retrieve target databases and tables in an end-to-end manner, taking into account structural relationships. (2) The generative retrieval paradigm makes it infeasible to perform semantic mapping of unseen schemas because they are not indexed into model parameters. In practice, labeling NL questions for massive database schemas is highly expensive. To this end, we propose a reverse schema-to-question generation paradigm to automatically synthesize training data and efficiently train the schema router. (3) How to incorporate routed schemas with LLMs to perform schema-aware SQL generation is critical to the quality of NL2SQL. We explore various prompt strategies to select and incorporate multiple candidate schemas for DBCopilot.
We use three standard NL2SQL datasets and two additional robustness datasets to evaluate DBCopilot. Experimental results verify the effectiveness of DBCopilot, with its copilot model outperforming retrieval-based baselines in schema routing by up to 19.88% in recall, and its schema-agnostic NL2SQL showing an improvement in execution accuracy of more than 7.35%. In summary, we demonstrate that DBCopilot is a scalable and effective solution for schema-agnostic NL2SQL, presenting a significant step forward in handling large-scale schemas of real-world scenarios.
Contributions.  Our contributions can be summarized as follows.
  • We highlight the limitations of existing NL2SQL methods in scaling from limited tables to massive databases.
  • We present DBCopilot, a framework that decouples schema-agnostic NL2SQL into schema routing and SQL generation via an LLM-copilot collaboration architecture.
  • We propose a relation-aware, end-to-end joint retrieval approach for schema routing, and also propose a reverse generation paradigm for automatic training data synthesis.
  • We conduct thorough experiments to verify the effectiveness of our solution in scaling NL2SQL to massive databases.
2. Preliminaries
2.1. Problem Definition
Table 1. Notation list used in this paper.
Symbol Description
$N$ Natural language question
$Q$ SQL query equivalent to question $N$
$\mathcal{D}$ Set of databases available, each denoted by $D$
$T$ Set of tables involved in query $Q$, with their properties
$S$ SQL query schema $\langle D,T\rangle$, where $T$ is the subset of tables belonging to database $D$
$\mathcal{T}$ Union of sets of tables from each database in $\mathcal{D}$
$\mathcal{S}$ Schema space $\mathcal{D}\times\mathcal{P}(\mathcal{T})$, where $\mathcal{P}(\mathcal{T})$ is the power set of $\mathcal{T}$
$\mathcal{G}$ Schema graph $\langle\mathcal{V},\mathcal{E}\rangle$, where $\mathcal{V},\mathcal{E}$ are graph nodes and edges
As mentioned above, we formulate the problem of natural language querying over massive databases and large numbers of tables with unknown SQL query schema (i.e., target database and tables) as schema-agnostic NL2SQL. Formally, as per the notation list detailed in Table 1, the problem of schema-agnostic NL2SQL is defined in Definition 2.1, and its two subtasks schema routing and SQL generation are defined in Definition 2.2 and Definition 2.3, respectively. Obviously, traditional NL2SQL with an NL question and its limited database schema as input is similar to the subtask SQL generation.
Definition 0 (Schema-Agnostic NL2SQL).
Given a natural language question $N$, schema-agnostic NL2SQL generates an equivalent SQL query $Q$ that can be executed against the appropriate database $D$ drawn from a collection of databases $\mathcal{D}$, without being provided with the intended SQL query schema $S=\langle D,T\rangle$.
Definition 0 (Schema Routing).
Schema routing is the process of identifying the SQL query schema $S=\langle D,T\rangle$ of a natural language question $N$ from the entire space of all possible schemas $\mathcal{S}=\mathcal{D}\times\mathcal{P}(\mathcal{T})$ within the specified set of databases $\mathcal{D}$ and tables $\mathcal{T}$.
Definition 0 (SQL Generation).
SQL generation is the process of generating a SQL query $Q$, given a natural language question $N$ and the intended SQL query schema $S=\langle D,T\rangle$.
Schema Linking vs. Schema Routing.  Schema linking (Katsogiannis-Meimarakis and Koutrika, 2023) is an optional substep of traditional NL2SQL methods. Given an NL question and a database schema, schema linking aligns NL terminologies, called “mentions”, with their corresponding schema elements (such as tables, columns, or database values). In this way, schema linking servers as a bridge between NL questions and database elements.
Schema linking and schema routing can be distinguished in two aspects. First, schema linking is a substep of SQL generation for accurate generation, while schema routing is a prerequisite of SQL generation to identify the unknown schema of a question. Second, the input for schema linking is the NL question along with the schema, whereas, the input for schema routing is the NL question and the output is its indexed or memorized schema.
2.2. Generative Retrieval
Traditional information retrieval (IR) techniques typically follow the “index-retrieve” paradigm. However, such a pipelined framework, which cannot be optimized together in an end-to-end manner, limits the final retrieval quality. To address this concern, generative retrieval (Tay et al., 2022; Wang et al., 2022) (a.k.a. DSI) is an emerging paradigm that trains a Seq2Seq pre-trained language model (PLM) to map queries directly to relevant document identifiers. The main idea behind it is to completely parameterize traditional “index-retrieve” pipelines within a single neural model, by modeling indexing and retrieval operations into the model parameters. Critical ongoing research in this paradigm focuses on defining better document identifiers (Sun et al., 2023; Zhang et al., 2023). To the best of our knowledge, this paper is the first to adapt this paradigm to jointly retrieve multiple elements with structural relationships.
3. Methodology
This section introduces DBCopilot, a schema-agnostic NL2SQL framework for effectively scaling NL querying to massive databases.
3.1. Overview
As shown in Figure 1, DBCopilot decouples natural language querying over massive databases into schema routing and SQL generation through LLM-Copilot collaboration. Specifically, given an NL question $N$, DBCopilot utilizes a copilot model as a schema router to identify the SQL query schema $S=\langle D,T\rangle$ associated with that question. After that, DBCopilot utilizes an LLM to translate the NL question $N$ with the routed schema $S$ into a corresponding SQL query $Q$. Example 3.1 provides an illustration of this process.
Example 0.
Given a question “Which singers held concerts in 2022?”, DBCopilot converts it into an equivalent SQL query through the following two phases.
1. Schema Routing identifies the relevant database and tables that contain the information required to answer the question:
[…]
2. SQL Generation prompts the LLM with the identified schema and question, and generates the SQL query:
SELECT s.name FROM singer_in_concert AS sic JOIN singer AS s
ON sic.singer_id = s.singer_id JOIN concert AS c
ON sic.concert_id = c.concert_id WHERE c.year = 2022
Figure 2. The training process of the schema router.
In practice, the schema space $\mathcal{S}=\mathcal{D}\times\mathcal{P}(\mathcal{T})$ is huge because it is the cartesian product of the database collections $\mathcal{D}$ and the power set of all tables $\mathcal{T}$. To narrow down the search space, we propose a relation-aware, end-to-end joint retrieval approach for schema routing. The training process is illustrated in Figure 2. Specifically, we first construct a schema graph to represent the relationships between databases and tables (§ 3.2). Based on this graph, we design a schema serialization algorithm to serialize a SQL query schema into a token sequence so that it can be generated by Seq2Seq DSI (§ 3.3). To address the challenge that DSI cannot generalize to unseen schemas, we further propose a reverse generation paradigm for training data synthesis (§ 3.4). These measures allow us to effectively train the schema router using serialized pseudo-instances ¡$N$, $S$¿ and apply inference with graph-based constrained decoding (§ 3.5). Finally, we explore various prompt strategies to select and incorporate multiple candidate schemas with LLMs (§ 3.6).
3.2. Schema Graph Construction
Unlike NL expressions, a valid SQL query must follow the structure of the database. Therefore, it is necessary to inform the schema router about the inherent relationships/constraints between schema elements. In general, the schema $S=\langle D,T\rangle$ involved in a valid SQL query must satisfy the following conditions. First, the set of tables $T$ must be the subset of tables belonging to database $D$, since a SQL query obviously cannot reference tables that do not exist in the target database. Second, the tables in $T$ should cross-reference at least one other table with a common column (e.g., foreign key), so that the SQL query can link them together via join or subquery clauses. In addition to explicit Primary-Foreign relations, two tables can also be linked by implicit Foreign-Foreign relations, i.e., they reference the same unique key column of another table.
Example 0.
Given a question “Which rivers run through the state with the largest city in the US?”, the corresponding SQL query “SELECT river_name FROM river WHERE traverse IN ( SELECT state_name FROM city WHERE population = ( SELECT MAX ( population ) FROM city ) )” uses only the city and river tables, both of which refer to the primary key of the state table.
To this end, we construct a directed graph $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$ to represent the underlying schema structure of all databases $\mathcal{D}$ and their tables $\mathcal{T}$. Concretely, the schema graph consists of three types of nodes $\mathcal{V}=\{\nu_{s}\}\cup\mathcal{D}\cup\mathcal{T}$, where $\nu_{s}$ is a special node denoting the set of all databases $\mathcal{D}$. Each edge $e\in\mathcal{E}$ in this graph represents a relation between two nodes, which can be divided into:
  • Inclusion relations: indicate whether a database is part of a database collection and whether a table belongs to a database.
  • Table relations: include explicit Primary-Foreign and implicit Foreign-Foreign relations between table nodes.
1
Input: A dictionary of all database schemas $databases$.
Output: The heterogeneous directed schema graph $\mathcal{G}$.
2 $\mathcal{G}\leftarrow$ an empty directed graph
3 for $database$ in $databases$ :
4       $links\leftarrow$ a dictionary with empty set as default value
5       $\mathcal{G}$.addEdge($\nu_{s}$, $database$)
6       for $table$ in $database$ :
7             $\mathcal{G}$.addEdge($database$, $table$)
8             for $column$ in $table$ :
9                   if $column$ contains a foreign key :
10                         $ltbl$, $lcol$leftarrow $column.{\mkern-2.0muf\mkern-2.0mu}oreign\_key$
11                         Add $ltbl$ and $table$ into set $links[ltbl.lcol]$
12
13
14
15
16
17
18      for $tables$ in $links$ :
19             for $tbl1$, $tbl2$ in $tables\times tables$ :
20                   if $tbl1$ is not equal to $tbl2$ :
21                         $\mathcal{G}$.addEdge($tbl1$, $tbl2$)
22
23
24
25
26
27
28
29
30return $\mathcal{G}$
Algorithm 1 Schema Graph Construction.
Algorithm 1 shows the procedure for constructing this three-tiered hierarchical schema graph. By capturing the structure of database schemas in graph $\mathcal{G}$, any valid SQL query schema $S$ is a trail (Wikipedia contributors, 2023b) on this graph. Each schema trail starts with node $\nu_{s}$, followed by a database node and some of its interconnected table nodes. The schema graph can efficiently narrow down the schema space $\mathcal{S}$ into trials on this graph starting from node $\nu_{s}$. As a result, the schema graph is essential for effectively representing and sampling SQL query schemas, serving as the foundation of DBCopilot.
3.3. Schema Serialization
Since we model schema routing as a Seq2Seq generative retrieval process, we need to design a schema serialization algorithm that converts a SQL query schema into a sequence of tokens.
Basic Serialization.  A trivial solution is to randomly order the set of tables as ad hoc retrieval (Chen et al., 2023), and identify the database based on the votes of the generated tables. However, this method loses sight of the inclusion and table relations between schema elements. The former results in a huge search space due to the lack of database constraints during table generation. The latter can result in missing tables during schema routing, especially junction tables.
Depth-First Search Serialization.  To address the above problem, we follow the constructed schema graph and design a depth-first search (DFS) serialization algorithm. As shown in Algorithm 2, we perform a DFS starting at node $\nu_{s}$ until all nodes of a SQL query schema have been visited and record the access order. We then concatenate the sequence of node names into a sequence of tokens. Formally, the depth-first search serialization can be defined as:
$\operatorname{Serialize}_{\pi}(S)\coloneqq\operatorname{Concat}\left(\left[\operatorname{DFS_{\pi}}(S,\mathcal{G})[i]\right]_{i=1}^{\left|\operatorname{DFS_{\pi}}(S,\mathcal{G})\right|}\right)$
where $\pi$ is a node iteration order and $\operatorname{DFS_{\pi}}(S,\mathcal{G})$ is the DFS traversal sequence of schema $S$ on schema graph $\mathcal{G}$ when iterating successor nodes under order $\pi$.
We introduce the iteration order parameter $\pi$ to determine the DFS sequence because a SQL query schema is not a completely ordered set and there may be multiple DFS sequences for a schema. We randomly select one DFS sequence at a time as the model target during training and assign different probabilities to different sequences originating from the same schema during inference.
1
Input: An ordered pair of a SQL query schema $S=\langle D,T\rangle$, the constructed schema graph $\mathcal{G}$, and a node iteration order $\pi$.
Output: The serialization of schema $S$.
2 $nodes$leftarrow $\{\nu_{s}\}\cup\{D\}\cup T$
$visited$leftarrow [][\;][ ]
  /* an empty list */
3 $stack$leftarrow $[\nu_{s}]$
4 while $stack$ is non-empty :
5       $node$leftarrow $stack$.pop()
6       $visited$.append($node$)
7       if set($visited$) is equal to $nodes$ :
             break
              /* cease the loop */
8
9
10      $successors$leftarrow [][\;][ ]
11       for $successor$ in $\mathcal{G}$[$node$] :
             /* Iterates over successor nodes according to node iteration order $\pi$/
12             if $successor$ in $nodes$ and not in $visited$ :
13                   $successors$.append($successor$)
14
15
16
17
18      $stack$.extend($successors$)
19
20
return $\operatorname{Concat}(visited[1:\operatorname{len}(visited)])$
  /* skip $\nu_{s}$/
Algorithm 2 Depth-First Search Serialization.
3.4. Training Data Synthesis
Based on the schema graph and the DFS serialization, we can learn the schema router on question-schema pairs ($N$, $S$) as described in § 3.5. However, the generative retrieval paradigm prevents the schema router from generalizing to unseen schemas because they are not injected into the model parameters. To address this challenge, we propose a labor-free training data synthesis method that automatically generates pseudo-instances via a reverse schema-to-question generation paradigm, as shown in Figure 2.
Specifically, we sample a sufficient number of valid schemas from the schema graph and generate a pseudo-question for each sampled schema. These schemas are sampled by performing finite-length random walks over the schema graph, and the database and tables traversed by a walk form a sampled schema. To generate an NL question for each sampled schema, we train a schema questioning model $\mathcal{M}_{\text{q}}$ in reverse based on the training sets of NL2SQL datasets. We use a SQL parser to extract the metadata (tables and columns) of SQL queries from the training pairs ($N$, $Q$). Unlike the schema router, which takes a condensed SQL query schema as output, the schema questioner accepts more detailed schema information (such as table columns) as input to generate rich and detailed questions. Given an illustration of the model input and output in Figure 3, the training loss of the schema questioning model $\mathcal{M}_{q}$ with parameters $\theta_{\text{q}}$ on $m$ extracted schema-question pairs ($S$, $N$) is:
$\mathcal{L}_{q}=-\sum_{i=1}^{m}\log P\left(N_{i}\mid S_{i};\theta_{\text{q}}\right)$
Pseudo-Question Generation Input: Ask a question that needs to be answered by combining the contents of all the tables, based on the word1 database schema provided below. - countrylanguage(language, countrycode, official) - country(code, lifeexpectancy) Output: What is the official language of the country with the highest life expectancy?
Figure 3. An illustration of pseudo-question generation.
3.5. Schema Router Training & Inference
Through the above measures, we can efficiently train the schema router to parameterize the semantic mapping of massive database schemas and effectively route NL questions to their target schemas via graph-based constrained decoding. Formally, the Seq2Seq schema router $\mathcal{M}_{r}$ takes a question sequence $N=\left[t_{1},t_{2},\ldots,t_{|N|}\right]$ as input and generates the serialized schema sequence $\bar{S}=\left[s_{1},s_{2},\ldots,s_{|\mathbf{s}|}\right]$. Suppose the synthetic data is $\{(N_{i},S_{i})\mid 1\leq i\leq n\}$, the training loss of schema router $\mathcal{M}_{r}$ with parameters $\theta_{r}$ is:
$\mathcal{L}_{r}=-\sum_{i}^{n}\sum_{\pi}\log P\left(\bar{S}_{i}\mid N_{i};\theta_{r}\right)$
After training the schema router, we further design a graph-based constrained decoding algorithm to guarantee the validity of generated schemas. As shown in Figure 4, at each step of autoregressive decoding, we maintain a dynamic prefix tree containing the names of accessible nodes from decoded schema elements. Unlike the DFS serialization, the accessible table nodes can only be the neighbors of generated tables, not the neighbors of the database, unless no tables have been generated yet. Therefore, the available tokens (part of the node name) for each decoding step can be obtained by searching the prefix tree, prefixed with the tokens generated after the last element separator. Besides, multiple candidate schema sequences can be generated to improve the coverage of the target schema. We thus utilize diverse beam search (Vijayakumar et al., 2016), which allows for generating a more diverse set of candidate sequences.
Figure 4. An example of graph-based constrained decoding.
3.6. SQL Generation
By decoupling schema-agnostic NL2SQL into schema routing and SQL generation, DBCopilot can easily integrate with orthogonal advanced LLM-based NL2SQL solutions, no matter whether in-context prompt engineering methods (Pourreza and Rafiei, 2023; Gao et al., 2023; Guo et al., 2023; Chang and Fosler-Lussier, 2023) or specific NL2SQL LLMs (Defog.ai, 2023; Station, 2023). However, since the schema router can generate multiple schemas from different databases for an NL question, we explore three prompt strategies to select and incorporate these different database schemas for LLMs:
  • Best Schema Prompting: instructs LLMs with the highest probability schema generated by the schema router.
  • Multiple Schema Prompting: instructs LLMs using multiple candidate schemas by simply concatenating the table schemas from different sequences in the prompt.
  • Multiple Schema COT Prompting: instructs LLMs using multiple candidate schemas through chain-of-thought (COT) reasoning (Wei et al., 2022). The LLM is first asked to select the most relevant candidate schema and then generates SQL queries.
4. Experiments
In this section, we conduct experiments on schema routing and schema-agnostic NL2SQL to verify the effectiveness of DBCopilot.
4.1. Experimental Setup
4.1.1. Datasets
We adapt existing NL2SQL datasets to evaluate schema routing and schema-agnostic NL2SQL.
  • Spider (Yu et al., 2018): Spider is a widely-used cross-domain NL2SQL dataset with 10,181 questions and 5,693 unique SQL queries across 138 domains and 200 databases.
  • Bird (Li et al., 2023b): Bird is a recent challenging NL2SQL dataset designed to emphasize the importance of database values, spotlighting the challenges presented by dirty database content, external knowledge requirements, and SQL efficiency.
  • Fiben (Sen et al., 2020): Fiben is a dataset designed to evaluate the processing of complex business intelligence queries, often involving nesting and aggregation, for NL2SQL systems. It consists of 300 NL questions, corresponding to 237 different complex SQL queries on a database with 152 tables.
  • Spidersyn (Gan et al., 2021): It is a human-curated dataset designed to evaluate the robustness of NL2SQL models to synonym substitution. It builds on Spider with NL questions modified by replacing schema-related words with real-world paraphrases.
  • Spiderreal (Deng et al., 2021): This is another robustness dataset created from the dev split of the Spider dataset by manually modifying the original questions to remove the explicit mention of column names while leaving the SQL queries unchanged.
Table 2. Statistics of datasets. Asterisks (*) indicate that the original development sets are used as test sets.
Dataset Size Schema
Train Dev. Test # DB # Tables # Cols
Spider 7000 1034* 166 876 4503
Bird 9427 1534* 80 597 4337
Fiben - - 279 1 152 374
Spidersyn 7000 1034 166 876 4503
Spiderreal 508 166 876 4503
Dataset Adaptation.  These datasets are designed to evaluate NL2SQL on a single database, where each question is annotated with its target database. To approach the schema-agnostic NL2SQL, we make several adjustments. First, we remove the single-database constraint and treat all databases in the collection as potential query targets. Next, we extract metadata (tables and columns) from each SQL query using a SQL parser (Mao, 2023) and exclude any queries that cannot be parsed. Finally, we combine the target database and extracted metadata as a SQL query schema, together with the original NL question and SQL query, to form each instance ($N$, $S$, $Q$). This results in three database collections (Spider, Bird, and Fiben), with instances divided into training and test sets as in the original setup. The robustness variants of Spider share the same database collection. Table 2 summarizes the dataset statistics after adaptation.
Table 3. Schema routing performance on standard test sets. The best results are in bold and the second best are underlined.
Method Spider Bird Fiben
Database Table Database Table Database Table
R@1 R@5 R@5 R@15 R@1 R@5 R@5 R@15 R@1 R@5 R@5 R@15
Zero-shot
BM25 70.12 91.49 86.49 93.87 57.11 90.16 68.32 82.81 98.92 98.92 33.34 38.62
SXFMR 62.57 89.46 80.42 92.39 66.49 88.20 67.56 83.05 100.0 100.0 28.21 46.47
LLM-enhanced
[…]
CRUSHSXFMR 63.25 91.97 82.21 93.86 70.14 92.44 70.58 85.07 100.0 100.0 34.06 50.78
Fine-tuned
BM25 68.28 91.78 86.60 93.87 60.63 90.48 69.87 83.34 98.92 98.92 32.67 38.30
[…]
4.1.2. Baselines
We compare DBCopilot to various sparse/dense retrieval techniques for scaling NL querying over massive databases, including zero-shot, LLM-enhanced, and fine-tuned methods:
  • BM25 (Wikipedia contributors, 2023a): BM25 is a well-known ranking function widely used in search systems to retrieve relevant documents.
  • SXFMR (Reimers and Gurevych, 2019): SXFMR is a technique for obtaining generic embedding models for dense retrieval by contrastive learning of Transformer(XFMR)-based PLMs on related text pairs.
  • CRUSH (Kothyari et al., 2023): CRUSH adopts a two-stage approach to retrieve schema elements for NL2SQL. It first instructs the LLM to generate a hallucinated schema, and then combines multiple retrieval results to obtain the actual schema.
  • DTR (Herzig et al., 2021): DTR is an approach that uses contrastive learning on labeled (question, table) pairs to develop table retrievers for open-domain question answering over tables.
4.1.3. Evaluation Metrics
For schema routing, we use Recall@k and mAP (mean average precision) to evaluate the relevance of retrieved databases and tables. Recall@k measures the fraction of relevant instances in the top-k retrieved candidates. mAP calculates the average precision over all queries, which reflects the overall retrieval performance. For SQL generation, traditional surface-level metrics are inappropriate for LLM-based approaches as they may produce SQL queries that differ greatly from the ground truth. Following recent studies (Gu et al., 2023b; Gao et al., 2023; Li et al., 2023b), we use EX (execution accuracy), which compares the results of generated SQL queries with corresponding ground truths. We also report the cost of LLM invocations.
4.1.4. Implementation Details
We implemented DBCopilot using Python 3.10, Pytorch 2.1, and the Transformers (Wolf et al., 2020) library.
Schema Routing.  We adopted T5-base (Raffel et al., 2020) as the backbone of the schema questioner and router. We set the batch size to $32$, the learning rate to $5\times 10^{-5}$, and the maximum number of training epochs to $20$ for all models. Model parameters were optimized using AdamW (Loshchilov and Hutter, 2019) and a linear learning rate schedule with no warm-up steps. We used the schema questioner trained on Spider and Bird to generate pseudo-questions for Fiben, which has no training data. For each database collection, we synthesized $10^{5}$ training instances.
SQL Generation.  We used gpt-3.5-turbo-16k-0613 to select the optimal schema and generated the SQL query with the OpenAI demo prompt (Gao et al., 2023). We set the temperature to 0 for reproducibility.
Baselines.  To maintain fairness, all baselines compared in this study relied on schema information only. We treated tables as retrieval targets, with the content of each table being the flat normalized names of the table and its columns. For each NL question, we retrieved the top 100 tables and determined database scores based on the average of their table scores. We used the standard Okapi BM25 with two adjustable parameters for BM25 and adopted the best performing all-mpnet-base-v2 for SXFMR. BM25 and DTR were fine-tuned on synthetic data, consistent with DBCopilot.
4.2. Schema Routing
We first compare the performance of schema routing, with the results shown in Table 3, Table 4, Figure 5, and the following findings.
Table 3 shows the database and table recall on regular test sets. We can see that DBCopilot consistently outperforms all types of baselines by a large margin. The database recall@1 of DBCopilot surpasses the second-best approach by up to 12.67% and 11.93% on the Spider and Bird datasets, respectively. This result verifies that DBCopilot can effectively identify the target database of an NL question. Furthermore, DBCopilot also holds an advantage in table recall, with a lead of up to 9.59%, which is crucial for providing complete tables for SQL generation. Even on the single database dataset Fiben, DBCopilot still achieves a 3.4% improvement. These results confirm that DBCopilot can effectively navigate NL questions to their target databases and tables over massive databases.
Finding 1. DBCopilot consistently outperforms all types of baselines by a large margin in schema routing.
Table 4. Schema routing performance on robustness test sets.
Method Spidersyn Spiderreal
Database Table Database Table
R@1 R@5 R@5 R@15 R@1 R@5 R@5 R@15
Zero-shot
BM25 29.11 58.03 45.30 58.87 52.95 81.1 70.87 80.64
[…]
LLM-enhanced
CRUSHBM25 35.88 65.18 46.80 64.06 58.46 86.42 75.38 89.14
[…]
Fine-tuned
BM25 30.66 57.83 46.12 59.65 50.00 81.30 70.77 80.87
[…]
DBCopilot 62.67 85.11 70.35 86.26 79.53 95.47 87.93 96.06
When querying over massive databases, data consumers without schema knowledge typically express NL questions that are highly inconsistent with database schemas. To evaluate the robustness in this situation, Table 4 shows the results of Spidersyn and Spiderreal, which replace explicitly mentioned schema elements in NL questions posed by NL2SQL annotation. We can see that the performance of all methods decreases with this diversity setting. Sparse retrieval baselines are significantly affected as they rely heavily on the vocabulary matching assumption, and the LLM-enhanced method also struggles with this challenge, despite being designed specifically for this case. Compared to these baselines, DBCopilot is the least affected and therefore maintains a larger advantage with 14.89% and 19.88% on database recall@1, as well as 2.83% and 6.94% on table recall@5. We believe this is because by injecting schema knowledge through schema-specific training, our method can better understand the semantics of database schemas and effectively learn the mapping between questions and schemas.
Finding 2. DBCopilot is more robust when natural language questions are highly inconsistent with database schemas.
Figure 5. (a) Table mAP w.r.t. different numbers of database tables. (b) Table recall@$k$ w.r.t. the number of retrieved tables.
DBCopilot model schema routing as a relation-aware, generative joint retrieval process, which differs from traditional baselines of independent retrieval. To assess the effect of relation modeling and joint retrieval, we evaluate the performance under different numbers of database tables and retrieved tables. As shown in Figure 5 (a), the mAP of DBCopilot is relatively stable as the number of database tables increases, while the independent retrieval baselines decrease significantly. The recall of DBCopilot is also consistently higher than the baselines as the number of retrieved tables increases from Figure 5 (b). These results verify the scalability of DBCopilot through relation modeling and generative routing.
Finding 3. Relation modeling and joint retrieval are essential for DBCopilot to perform effective schema routing.
4.3. Schema-Agnostic NL2SQL
We then compare the schema-agnostic NL2SQL performance under gold schemas, different routing methods, and prompt strategies.
Oracle Test.  We conduct oracle tests to access the upper bound of schema-agnostic NL2SQL, by providing gold schemas at various small scales. Specifically, we gradually move from 5 database schemas to the gold database, then to gold tables only, or even to gold tables with only gold columns. Experimental results in Table 5 show that as the number of schema elements increases, the EX of LLM-generated SQL queries continuously decreases, while the inference cost keeps increasing. This suggests that LLMs, like PLMs (Li et al., 2023c), are also affected by extraneous schema elements in SQL generation. Predictably, the situation gets worse as the scale of the schema increases. As a result, it is highly inefficient and ineffective to expose massive database schemas directly to LLMs.
Table 5. Schema-agnostic NL2SQL performance with gold (G.) schemas, different routing methods, and prompt strategies.
Routing Method Provided Schema Prompt Strategy Spider Bird Spidersyn
EX Cost EX Cost EX Cost
G. Column Basic 79.69 0.40 35.98 0.75 76.60 0.40
G. Table Basic 77.08 0.52 29.86 1.44 67.21 0.52
G. Database Basic 74.47 0.81 25.75 2.70 63.73 0.81
5 Database Basic 67.89 2.98 19.10 8.80 49.90 2.99
CRUSHBM25 Top 1 Basic 56.77 1.30 13.69 2.67 20.79 1.23
DTR Top 1 Basic 45.16 0.66 17.60 2.12 26.98 0.62
DBCopilot Top 1 Basic 64.12 0.69 21.97 2.35 37.81 0.70
DBCopilot Top 5 Basic 61.51 2.28 19.56 7.18 34.53 2.27
DBCopilot Top 5 COT 57.83 5.33 19.17 16.49 33.95 5.29
[…]
Finding 4. Schema routing is necessary for LLM-based schema-agnostic NL2SQL for both efficiency and effectiveness.
Impact of Improved Recall.  We perform schema-agnostic NL2SQL to verify the impact of improved recall on the accuracy of generated SQL queries. Without loss of generality, we compare DBCopilot to CRUSHBM25 (sparse) and DTR (dense), which achieve suboptimal routing performance on Spider and Bird, respectively. As shown in Table 5, DBCopilot outperforms these two baselines by up to 10.83% EX. The LLM inference cost of DBCopilot is slightly higher, possibly due to successfully retrieving more required tables. We can also confirm the correlation between routing recall and EX by comparing these two baselines. CRUSHBM25 outperforms DTR in recall and EX on Spider, while DTR achieves higher EX on the other two datasets due to its superior recall. Therefore, the success of schema routing is essential for schema-agnostic NL2SQL.
Finding 5. DBCopilot outperforms retrieval baselines in schema-agnostic NL2SQL due to its superior schema routing.
Exploration of Prompt Strategies.  By generating multiple candidate schemas, DBCopilot can effectively increase the recall of the target schema. We explore three prompt strategies for leveraging multiple schemas. Interestingly, SQL generation with the top 1 schema performs best, even though the top 5 candidates have 22.44% more database recall on Spidersyn than the top 1. This may be due to confusion caused by similar candidate schemas, thus requiring more database/human information to disambiguate. By introducing human-in-the-loop (HITL) to select the desired schema from the top 5 candidates, DBCopilot comes close to the ideal situation of providing the gold database schema, but with lower inference costs. This highlights the ability of DBCopilot to exclude extraneous tables and its promise for schema-agnostic NL2SQL.
Finding 6. Using the best schema prompting strategy for SQL generation is optimal unless there is human interaction.
4.4. Ablation Study
Table 6. Performance loss in ablation studies. “BS”, “OD”, “CD”, “DB” stands for basic serialization, original NL2SQL training data, constrained decoding, and diverse beam search.
Spider Bird
Database Table Database Table
R@1 R@5 R@5 R@15 R@1 R@5 R@5 R@15
DBCopilot 85.01 96.42 91.63 97.51 88.92 97.13 85.83 94.59
   w/ BS -3.20 -2.90 -4.09 -5.75 -3.39 -1.63 -6.43 -5.99
   w/ OD -84.72 -82.40 -88.76 -79.55 -88.92 -82.27 -85.73 -88.81
   w/o CD -0.39 -0.19 -4.15 -2.51 +0.12 -0.45 -2.81 -1.93
   w/o DB +0.48 -3.00 +0.64 -2.88 -0.52 -4.89 -3.69 -8.87
We finally perform ablation studies on DBCopilot to validate the contribution of each module, as summarized in Table 6.
DFS Serialization.  DFS serialization is essential for schema routers to learn the underlying relations of databases and tables. To demonstrate this, we replace the DFS serialization with the unordered basic serialization of tables, similar to DSI for multiple documents. As shown in Table 6, this results in a 1% to 6% reduction in recall. DFS serialization effectively reduces the search space, whether by first generating a database and then generating its tables, or by continuing to generate based on existing tables. This graph-based serialization also makes constrained decoding possible.
Data Synthesis.  Generative retrieval is difficult to generalize to unseen documents during training. As we can see, routers trained on the original training instances ($N$, $S$) extracted from NL2SQL datasets are almost completely incapable of routing on test schemas, even when constrained decoding is used. Data synthesis by sampling schemas and generating pseudo-questions can effectively solve this challenge. Figure 6 shows how the performance varies with the amount of synthetic data. Database and table recalls increase rapidly at first as more data is synthesized, but level off after a certain amount. How to improve the quality of synthetic data and reduce the need for it may be a future direction.
Decoding Strategy.  Decoding optimizations (i.e., constrained decoding and group beam search) are beneficial for generative schema routing during inference. On the one hand, constrained decoding is useful for generating accurate table identifiers, since the drop from removing it is mainly in table recalls. On the other hand, group beam search helps to generate multiple diverse schemas because database recall@5 decreases significantly without it.
Finding 7. DFS serialization, data synthesis, and decoding optimizations contribute to DBCopilot in different ways.
5. Related Work
5.1. Natural Language to SQL
NL2SQL is an essential part of the natural language interface. Most research has approached this task along the lines of semantic parsing (Kamath and Das, 2019; Li et al., 2020). Early NL2SQL parsing systems adopted symbolic approaches that relied heavily on rule or grammar engineering, making them difficult to apply widely (Zelle and Mooney, 1996; Woods, 1973). In recent years, with the advances in deep learning, especially PLMs, end-to-end NL2SQL approaches that using neural generation models have become the mainstream (Li et al., 2023c; Wang et al., 2020; Gu et al., 2023a; Li et al., 2023a; Yu et al., 2021). Some efforts have been made to improve encoding and decoding, such as relation-aware encoding (Wang et al., 2020; Li et al., 2023a) and grammar-based decoding (Xu et al., 2017; Scholak et al., 2021). Some other efforts have focused on generating accurate SQL queries through schema linking (Li et al., 2023c; Wang et al., 2020) or skeleton generation (Gu et al., 2023a, b).
LLMs have emerged as a new paradigm for NL2SQL (Rajkumar et al., 2022; Liu et al., 2023a; Gao et al., 2023). Unlike previous work, the core of the LLM-based NL2SQL solution is prompt engineering, i.e., prompting LLM to generate correct and efficient SQL queries. Several efforts have been made to select appropriate in-context learning examples through question classification (Pourreza and Rafiei, 2023), skeleton retrieval (Guo et al., 2023), and in-domain example synthesis (Chang and Fosler-Lussier, 2023). Research has also been done on decoupling zero-shot NL2SQL into PLM-based sketch generation and LLM-based SQL completion (Gu et al., 2023b). These studies are orthogonal to our work and can be easily integrated to optimize DBCopilot SQL generation.
5.2. Table Discovery
Table discovery techniques are critical to data management, allowing users to explore and gain insights from data warehouses or lakes (Fan et al., 2023a). A common practice for table discovery is query-driven discovery (Zhang and Ives, 2020; Nargesian et al., 2019), where the query is typically in the form of keywords, tables, or NL questions. For keyword search, data discovery systems find relevant tables by searching for topic terms in the metadata (Cafarella et al., 2009; Pimplikar and Sarawagi, 2012). For table search, tables are retrieved to augment a given table as the query (Sarma et al., 2012). Joinable table search (Dong et al., 2021; Zhu et al., 2019) and unionable table search (Fan et al., 2023b; Khatiwada et al., 2023) aim to augment a query table with additional attributes and new tuples, respectively. For NL question search, open domain question answering retrieves independent tables to answer the question with table-based models (Chen et al., 2021; Herzig et al., 2021). Although dense retrieval methods based on contrastive learning have achieved advanced performance in the above tasks (Dong et al., 2021; Fan et al., 2023b; Herzig et al., 2021), they are difficult to consider the structural relationships between tables and retrieve them jointly. Alternatively, data navigation allows users to navigate interactively through the tables organized hierarchically (Nargesian et al., 2020) or relationally (Fernandez et al., 2018). Our work can be seen as automated data navigation, where we employ DSI to model the query-based navigation in an end-to-end manner.
Figure 6. Database recall@1 and table recall@5 w.r.t. the number of (#) synthetic training data.
6. Conclusion
In this paper, we present DBCopilot, a framework for effectively scaling natural language querying to massive databases. DBCopilot decouples schema-agnostic NL2SQL into schema routing and SQL generation through LLM-Copilot collaboration. We present an approach to effectively navigate the target database and tables for an NL question in a relation-aware, end-to-end manner. We also propose a training data synthesis paradigm to adapt routers over massive databases without manual intervention, and explore various prompt strategies to incorporate multiple candidate schemas. Extensive experiments verify the effectiveness of our solution. In summary, DBCopilot pushes the boundaries of NL2SQL, enabling researchers to better strategize their quest for data accessibility.
Appendix A Supplementary Experiments
A.1. Case Study
[…]
DPR $\langle\text{world},\{\text{countrylanguage}\}\rangle$
[…]
Figure 7. Successful case where DBCopilot routes the correct schema but baselines fail.
[…]
We provide successful cases to illustrate how DBCopilot improves schema routing through joint retrieval. As the example in Figure 7 shows, DBCopilot is the only approach that accurately routes the appropriate schema for the given natural language question. Sparse retrieval methods (BM25 and CRUSHBM25) rely heavily on lexical matching while ignoring the semantic matching between the question and the schemas. However, untrained dense retrieval methods (SXFMR and CRUSHSXFMR) also fail to capture the underlying semantics of tables and retrieve the wrong databases. Fine-tuning based approaches identify the correct database where DPR misses the country table linked to the countrylanguage table, but relation-aware DBCopilot does not.
[…]
A.2. Efficiency & Resource Consumption
We also measure the efficiency and resource consumption of each schema routing method. Table 7 shows the number of queries per second (QPS), the duration of build time, the amount of disk space utilized, and the consumption of system and GPU memory. LLM inference and complex post-processing are the main reasons for the slowness of CRUSH approaches. DBCopilot, based on an encoder-decoder architecture, consumes more disk space and GPU memory than dense retrieval approaches (SXFMR, DPR) based on an encoder-only architecture. Despite the slightly higher resource consumption for training and inference in generative retrieval, we believe and have observed that the growing popularity of generative AI is leading to an increasing number of optimizations being proposed.
Table 7. Efficiency & resource consumption of different methods. Build time includes the time for model training and index construction. We calculate queries per second (QPS) and memory consumption with a query batch size of 64.
Method QPS Build (s) Disk (MB) RAM (MB) GPU (MB)
BM25 1677.1 10 36 800 -
SXFMR 49.0 48 420 2,230 3742
CRUSHBM25 0.9 10 36 1104 -
CRUSHSXFMR 0.8 48 420 2502 3742
BM25 1677.1 6972 36 800 -
DPR 49.0 5161 420 2,230 3742
DBCopilot 43.0 13678 852 656 11807
A.3. Oracle Test with In-Context Learning
Table 8. Schema-Aware NL2SQL with In-Context Learning
Provided Schema Prompt Strategy Spider Bird Spidersyn
EX Cost EX Cost EX Cost
G. Column Basic 77.95 3.52 35.98 11.58 76.69 3.43
G. Table Basic 75.82 3.61 28.49 12.10 69.34 3.52
G. Database Basic 75.73 3.83 25.75 13.08 63.64 3.74
5 Database Basic 68.38 5.57 21.45 17.79 50.48 5.47
Table 8 shows the results of the oracle test by providing few-shot (5-shot) examples. As we can see, while this improves the accuracy of NL2SQL somewhat, it does not compensate for the performance degradation caused by extraneous schema elements. This makes schema linking still important in LLM-based NL2SQL. In the future, we will investigate how to better disambiguate and link schema elements in the context of schema-agnostic NL questions.